OpenStudy (decarr432):

7^16 7^-18 ---- = ------ 7^12 ? The missing denominator is? A.7^(-6) B.7^(-14) C.7^(-22) D.7^22

OpenStudy (decarr432):

Any one have a clue?

OpenStudy (hwyl):

laws of exponent involving fractions \(\huge \frac{a^m}{a^n} = a^{m-n} \)

OpenStudy (decarr432):

huh?

OpenStudy (ikram002p):

do you know cross multiply ?? @Decarr432

OpenStudy (decarr432):

I'm taking a pretest on things iv'e never been taught and yes i know cross multiplication

OpenStudy (hwyl):

figure out the left hand side of the equation using the laws of exponent first then the right side of the equation will flow

OpenStudy (decarr432):

Okay one minute please

OpenStudy (ikram002p):

know what @hwyl idea is better :P

OpenStudy (hwyl):

do not use a calculator, just use the laws of exponent that I laid out

OpenStudy (decarr432):

So solve the exponents first right?

OpenStudy (decarr432):

or would I just 7^4?

OpenStudy (hwyl):

correct

OpenStudy (hwyl):

now that we figured the left hand side, let us look at the right hand side

OpenStudy (hwyl):

let us just focus on the exponent since we know that all of the base are 7 anyway. 4 = -18 - (?)